• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name yez

    Log entry time 13:09:17 on February 23, 2011

    Entry number 348271

    keyword=Several unsolved issues in HRSs (in general work fine)

    All indexes start from 0
    PRL: I checked all PMT signals which indicate pedestal-only or big
    noisy(only PRL1.20) and saw all of them in the scope, which means that
    they could be only problems of bad cables or ADC channels.
    1, PRL1:17,18,19,32 only have pedstal, PRL1.20 has broad pedestal ->I had
    no time to move the cable to other channels but will do it tomorrow 
    2, PRL2 15,16,32,33: We moved this cable from slot.25:0~15 to 16~31, but
    data taken after moving still show the problems. We will try to replace
    their cable or connection.
    The good news is that all signals are there so we just need to either
    change cables or switch ADC channels to bring all PRL back to work.
    1,Shower 41,73,75,77 have pedestal only, 40 has low ADC values and low
    All those PMTs are far from the center of detectors and their signals
    suppose to be small even they function normal, so we might not need to
    worry too much about them if we really can not fix them in time.
    All other detectors in HRSs work fine (so far:>).

    A copy of this log entry has been emailed to: doug@jlab.org, solvigno@jlab.org,kalyan@jlab.org,vasulk@jlab.org